1. A mass of 250 g stretches a spring 1.568 cm. If the mass is set in motion from its equilibrium position with a downward velocity of 20 cms, and if there is no damping, determine the position u of the mass at any time t.

Enclose arguments of functions in parentheses. For example, sin(2x).

a. Assume g = 9.8 m/s^2. Enter ans exact answer.
u(t) = ______ m

b. When does the mass first return to its equilibrium position?
Enter an exact answer.

t = ______s
1
Expert's answer
2020-07-06T20:21:24-0400

"m=0.25" kg

"L=0.01568" m

"u(0)=0, \ u^{\prime}(0)=0.2" m/s, where "u(t)" is position of the mass at any time "t."


"k=\frac{mg}{L}=\frac{0.25\times 9.8}{0.01568}=156.25" N/m


Now we can write the equation of the system: "mu^{\prime \prime }+ku =0"

"0.25u^{\prime \prime }+156.25u =0"

"u^{\prime \prime }+625u =0"

The characteristic equation: "r^2+625=0, \ \ r=\pm25i" .

The general solution is "\ u(t)=C_1\sin(25t)+C_2\cos(25t)"

"u(0)=0=C_1\times 0+C_2\times 1=C_2" , "C_2=0"

"u^{\prime }(0)=0.2=25C_1" , "C_1=0.008"

So, "u(t)=0.008\sin (25t)" m.

Equilibrium position: "u(t)=0, \ \ 0.008\sin(25t)=0, \ \ \sin(25t)=0, \ \ 25t=\pi n, \ n\in \mathbb{Z}" .

The mass first return to its equilibrium position when "n=1" or "t=\pi/25."


Answer:

a. "u(t)=0.008\sin (25t)" m.

b. "t=\pi/25" s.